# Evidence of Energy Injection in the Short and Distant GRB 250221A

### v4

@@ 5 Origin of Optical Rebrightening
Figure 1 shows evidence for a change in the fading evolution and a rebrightening at $T>0.6$ d. In this section, we explore and discuss possible scenarios that could account for this. Given the ambiguity surrounding the origin of the event, we first consider whether the observed excess can be explained by a supernova associated with the death of a massive star (Section 5.1), then we turn to scenarios related to the host galaxies and the surrounding environment, including the possibility of gravitational microlensing (Section 5.2) or density variations in the CBM (Section 5.3). Finally, we investigate whether the excess could be produced by a refreshed shock within the main jet of GRB 250221A (Section 5.4).
Scenarios involving kilonova origin for the optical excess can be ruled out. An AT2017gfo-like (Abbott et al., 2017) event placed at $z=0.768$ would appear at $r\sim 28.6$ mag, far below our detection limits. Even the brightest magnetar-powered kilonova models are undetectable at this distance (see e.g. Metzger et al., 2010; Metzger & Berger, 2012). Similarly, a fast blue optical transient (FBOT) (see e.g. Perley et al., 2019, 2021) origin is very unlikely. While these transients reach high luminosities and rise rapidly, the observed negative spectral slopes ($\beta<0$) of the afterglow of GRB 250221A are inconsistent with the expected blue continua of FBOTs ($\beta>0$). Thus, neither a kilonova nor an FBOT provides a plausible explanation for the excess emission.
@@ 5.1 Supernova
Broad-lined Type Ic SNe, have been associated with LGRBs through both temporal and spatial coincidence, as well as spectroscopic confirmation (see e.g. Galama et al., 1998; Hjorth et al., 2003; Pian & Mazzali, 2006; Mirabal et al., 2006; Becerra et al., 2017). These SNe usually present peak absolute magnitudes of $M_{V}\approx-19$ to 20-20 around 10 to 15 days after the burst in the rest frame (e.g. Cano et al., 2017). This timescale is determined by the radioactive decay of 56Ni and 56Co.
At the luminosity distance of $d=4.9$ Gpc, standard type Ibc supernovae (SN 1998bw-like; Galama et al., 1999) would exhibit apparent peak magnitudes in the range of $\sim~24.5.0-25.5$, making them several orders of magnitude fainter than the late-time component observed in GRB 250221A. On the other hand, a very luminous supernova associated with a magnetar such as SN2011kl could be brighter by about one magnitude Greiner et al. (2015) and therefore comparable with the values exhibited for GRB 250221A in the rebrightening.
@@ 5.2 Gravitational Microlensing
It is tempting to associate the rebrightening with a microlensing event, assuming that G2 is the host and G1 is a galaxy on the line-of-sight. In the following, we briefly describe the formalism for GRB microlensing first introduced by Loeb & Perna (1998). As discussed in section 4.2, the temporal evolution of the spherical blast wave is described by the Blandford & McKee (1976) self-similar solution, with $\Gamma\propto R^{-3/2}$, after it has been decelerated by its interaction with the ISM.
The afterglow image on the plane of the sky is predicted to appear as a ring with a narrow fractional width, $W$, that depends on the observed frequency and whose radius grows with apparent time $T$ as a power law, with $R_{s}(T)\propto R/\Gamma\propto R^{5/2}\propto T^{5/8}$ where $R\propto\Gamma^{2}T\propto T^{1/4}$. In the optical band, $W$ is predicted to be of the order of a few percent (Loeb & Perna, 1998; Granot et al., 1999). After about a day after the trigger, the radius of the ring, $R_{s}(T)\approx 1.1\times 10^{17}~E_{\rm k,iso,52}^{1/8}~n_{1}^{-1/8}T_{\rm day}^{5/8}~{\rm cm}$ (Piran, 1999), becomes comparable to the Einstein radius of a solar mass lens, $r_{E}=\sqrt{(4GM/c^{2})(D_{L}D_{LS}/D_{S})}$ where $D_{L}$, $D_{LS}$, and $D_{S}$ are the lens-observer, lens-source and source-observer angular distances, respectively, $M$ is the mass of the lens.
The magnification of a uniform ring of fractional width $W$ and impact factor $b$ is given by Loeb & Perna 1998,
@@
So, assuming that the lens is in G1 and the GRB is in G2, we get $D_{L}\approx 1037$ Mpc, $D_{LS}\approx 783$ Mpc, and $D_{S}\approx 1571$ Mpc, respectively (Planck Collaboration et al., 2020) and hence, $r_{E}\approx 3.1\times 10^{16}(M/M_{\odot})^{1/2}~{\rm cm}$. Since the radius of the ring depends on the kinetic energy and the density of the medium, we adopt here a range of values around the values shown in Figure 8. We find that for the following ranges $1\lesssim n_{1}\lesssim 5$ and $0.1\lesssim E_{\rm k,iso,52}\lesssim 10$, the timescale for the excess demands $b<1$ and $M\sim 50~M_{\odot}$. The width of the excess demands a fractional width of the ring $W\sim 0.1$. However, this modelling does not allow accounting for the amplitude of the excess, as can be seen in Figure 9. We note that these microlensing calculations are done considering a uniform ring for the afterglow image (Loeb & Perna, 1998).
@@ 5.3 ISM Local Density Enhancement
As we describe in Section  4.2, the fireball model describes the afterglow emission as a smooth light curve that can be modelled with power-law segments. However, in reality, light curve variability is observed in several GRBs (Nakar et al., 2003; de Ugarte Postigo et al., 2005), and can be attributed to an increase in external density that the fireball interacts with, experiencing additional deceleration (Nakar & Granot, 2007).
@@
In contrast, for SGRBs, and likely GRB 250221A, this explanation does not apply. Their compact-object progenitors merge long after both stars have exploded as supernovae, leaving behind a CBM free of stellar-wind disturbances. Thus, any external density variations must stem from inhomogeneities in the ISM (Nakar & Granot, 2007).
Nevertheless, Nakar & Granot (2007) demonstrated that density fluctuations in the external medium have a minimal impact on the forward shock emission as the dynamics of the relativistic blast wave tend to suppress any resulting re-brightening. Furthermore, van Eerten et al. (2009) performed hydrodynamical simulations of a relativistic blast wave interacting with a wind termination shock, finding no significant variability in the optical light curve, even in the presence of large density contrasts. Therefore, these results favour scenarios involving refreshed shocks or multiple ejection episodes as more plausible explanations for the observed variability.
@@ 5.4.1 Reverse Shock Emission from Two-Shell Collision
To explain the rebrightening of the afterglow emission around starting from $T_{0}\sim 1$ day (hereafter $T_{\rm bright}$) , we first consider a scenario where a second (inner) shell of mass emitted by the central engine, with some delay of $t_{\rm em}$ in the engine frame, catches up from behind with the decelerating first (outer) shell and refreshes the shock by injecting more energy (e.g. Kumar & Piran, 2000; Sari & Mészáros, 2000; Zhang & Mészáros, 2002; Vlasis et al., 2011). We develop the model of two spherical shell collision in Appendix (A), where we calculate characteristic radii, timescales, and dynamics of the system (also see, e.g., Anderson et al., 2025). As the two shells collide (see Figure 12) a forward shock (FS) propagates into the outer shell and a reverse shock (RS) propagates into the inner shell. The shock jump conditions at the two shocks are different, as the outer shell is already relativistically hot while the inner shell is cold. The main parameters in the problem include the ratio of the kinetic energies of the two shells ($E_{\rm k,iso,6}/E_{\rm k,iso,3}$), the ratio of their initial bulk Lorentz factors ($\Gamma_{0,3}/\Gamma_{0,6}$), and the shock microphysical parameters at the two shocks.
In the case of a binary merger, the two shells must be emitted with negligible delay, such that $t_{\rm em}\ll T_{\rm 90}$, and therefore the inner shell must have $\Gamma_{0,6}<\Gamma_{0,3}$. This condition also guarantees that the inner shell must be spreading radially before the collision, i.e. it is in the thin-shell regime. Also, in general, if $t_{\rm em}\ll T_{\rm bright}$, then the ratio of the initial bulk Lorentz factors is fixed (see Equation 16), such that
$\frac{\Gamma_{0,3}}{\Gamma_{0,6}}$ $\approx\frac{[4(T_{\rm bright}/T_{\rm dec})]^{3/8}}{2}$
@@
Since $\Gamma_{0,3}\sim 400$ is obtained from the MCMC fits, this yields $\Gamma_{0,6}\sim 8$. Figure 13 shows the dynamical evolution of the different shocked and unshocked regions, where we find that the RS remains non-relativistic as it crosses the inner shell, as expected for a thin shell.
As shown in Zhang & Mészáros (2002), the RS emission typically dominates over that coming from the other two shocked regions before it declines rapidly. The shock jump conditions in Equation 19 yield the energy and number density of the shocked medium, which we then use to calculate the characteristic synchrotron frequencies and the peak flux (e.g. Sari et al., 1998; Gill & Granot, 2018). In calculating the observed emission, we simplify to obtain radiation emitted only along the LOS without performing EATS integration, with photon arrival time given by $T=(1+z)R/4\Gamma^{2}c$.
@@
The two shell collision scenario, as described here, demands a significantly large amount of energy to be residing in the inner shell. In this case $E_{\rm k,iso,6}/E_{\rm k,iso,3}\simeq 20$, which for $\Gamma_{0,6}\sim 8$ implies a large density in the inner shell compared to the outer one. Since the RS is non-relativistic, large densities are needed to obtain a bright emission. When such a large amount of energy is injected into the blast wave, it should also lead to very bright FS emission (not calculated in this model) after the RS emission fades, which would significantly overproduce the emission during the decaying phase of the rebrightening episode. Therefore, this model is disfavoured by the observations.
@@ 5.4.2 Energy Injection into the Forward Shock of a Jet
Next we consider an alternative model with continuous energy injection into the blast wave produced by a jet. Two different ways of doing so have been discussed in the literature for a spherical flow, where (i) the ejecta comprises a radial velocity stratification, with progressively slower moving inner shells trailing behind the faster outer shell (Rees & Mészáros, 1998; Sari & Mészáros, 2000), or (ii) a rapidly spinning central engine, e.g. a millisecond magnetar, continuously injects energy into the blast wave via a magneto-hydrodynamical (MHD) wind as it spins down (Dai & Lu, 1998a, b; Zhang & Mészáros, 2001, 2002). In both scenarios, the afterglow light curve shows an achromatic bump or rebrightening as the energy is injected, with the post-injection light curve following the same temporal decay as that of the pre-injection one in the absence of spectral break passage. Both scenarios have been used in many earlier works (e.g., Laskar et al., 2018; Schroeder et al., 2024; de Wet et al., 2024; Schroeder et al., 2025) to explain rebrightening features in afterglow light curves, with the assumption that the trailing ejecta in model (i) only catches up with the faster moving blast wave at the time of the rebrightening.
@@
so that the energy of the blast wave grows as $E(R)\propto R^{1+\ell}$. From energy conservation in the blast wave (Blandford & McKee, 1976; Chiang & Dermer, 1999; Huang et al., 1999; Zhang & Mészáros, 2002), we get
@@
To solve for the dynamical evolution of the spherical blast wave with energy injection, we use the numerical code of Gill & Granot (2023) that produces EATS-integrated afterglow emission. Figure 11 shows the afterglow light curve where we take the best-fit solution from Figure 8 and inject $E_{\rm inj}\simeq 4.5E_{\rm k,iso}$ into the blast wave with $\ell=0.6$. During the energy injection phase, the light curves for the radio ($\nu_{m}<\nu_{R}<\nu_{a}<\nu_{c}$), optical ($\nu_{a}<\nu_{O}<\nu_{c}$), and X-rays ($\nu_{c}<\nu_{X}$) show distinct temporal trends since all three are located on different power-law segments of the synchrotron spectrum. Here $\nu_{m}$ and $\nu_{c}$ are the characteristic injection and cooling-break frequencies, as defined earlier, and $\nu_{a}$ is the self-absorption frequency. Since the radio emission is self-absorbed, the rebrightening shows a temporal lag of $\Delta T\sim 4$ days with respect to that seen in optical and X-rays.
To explain the late-time radio and X-ray emission, a jet break in the light curve at $T\sim 10^{5}$ s is required, with a jet half-opening angle of $\theta_{j}=11.5^{\circ}$. An on-axis ($\theta_{\rm obs}=0$) observer would see a jet break in the afterglow light curve when the angular size of the beaming cone around the line-of-sight exceeds the opening angle of the jet, i.e. when $1/\Gamma=\theta_{j}$, which would occur at the time $T_{j}=(\Gamma_{0}\theta_{j})^{8/3}T_{\rm dec}\simeq 2\times 10^{5}$ s for $\Gamma_{0}=400$ and $T_{\rm dec}\simeq 2$ s. The true jet kinetic energy in this scenario can be estimated from the total isotropic-equivalent energy, with $E_{k}\simeq(\theta_{j}^{2}/2)E_{\rm k,iso,tot}=4.4\times 10^{51}$ erg where $E_{\rm k,iso,tot}\simeq 5.5E_{\rm k,iso}$, and $E_{\rm k,iso}$ is the energy of the fastest moving material.
@@
In the first scenario, where the ejecta has a radial gradient in velocity, so that the mass above a given bulk-$\Gamma$ scales as $M(>\Gamma)\propto\Gamma^{-s}$ (Rees & Mészáros, 1998; Sari & Mészáros, 2000), the energy in the blast wave grows as $E(>\Gamma)\propto\Gamma^{1-s}\propto T^{-3(1-s)\over(7+s)}$ after the fastest moving material decelerates. The scaling of the blast wave energy with $T$ in this model corresponds to $s=(4+\ell)/(2-\ell)\simeq 2.2$. Since $s>1$, both the mass and energy in the ejecta are dominated by the slower moving material. In the second scenario, energy is injected by the spinning down magnetar with power $L(T)\propto T^{q}$, where the blast wave energy grows as $E\propto T^{1+q}$ (Zhang & Mészáros, 2001). Comparison of this scaling with our model gives $q=2(\ell-1)/(3-\ell)=-0.62$.

### v5

@@ 5 Origin of Optical Rebrightening
Figure 1 shows evidence for a change in the fading evolution and a rebrightening at $T>0.6$ d.
Late-time rebrightenings have been reported in several long GRBs: GRB 100814A (Nardini et al., 2014) and the ultra-long GRB 111209A (Kann et al., 2019; Gendre et al., 2013; Greiner et al., 2015). In addition, the Einstein Probe mission (Yuan et al., 2025) has recently identified rebrightening events without associated GRBs, for example EP241021a (Gianfagna et al., 2025; Yadav et al., 2025) and EP240414a (van Dalen et al., 2025), likely representing off-axis explosions. While these may constitute a different phenomenological class, they demonstrate that such late-time rises are not unprecedented among long-duration or SN-associated transients.
In this section, we explore and discuss possible scenarios that could account for the excess in GRB 250221A. Given the ambiguity surrounding the origin of the event, we first consider whether the observed excess can be explained by a kilonova, a fast blue optical transient, or a supernova associated with the death of a massive star (Section 5.3), then we turn to scenarios related to the host galaxies and the surrounding environment, including the possibility of gravitational microlensing (Section 5.4) or density variations in the CBM (Section 5.5). Finally, we investigate whether the excess could be produced by a refreshed shock within the main jet of GRB 250221A (Section 5.6).
5.1 Kilonova
Scenarios involving kilonova origin for the optical excess can be ruled out. An AT2017gfo-like (Abbott et al., 2017) event placed at $z=0.768$ would appear at $r\sim 28.6$ mag, far below our detection limits. Even the brightest magnetar-powered kilonova models are undetectable at this distance (see e.g. Metzger et al., 2010; Metzger and Berger, 2012).
5.2 FBOT
Similarly, a fast blue optical transient (FBOT) (see e.g. Perley et al., 2019, 2021) origin is very unlikely. While these transients reach high luminosities and rise rapidly, the observed negative spectral slopes ($\beta<0$) of the afterglow of GRB 250221A are inconsistent with the expected blue continua of FBOTs ($\beta>0$). Thus, neither a kilonova nor an FBOT provides a plausible explanation for the excess emission.
@@ 5.3 Supernova
Broad-lined Type Ic SNe, have been associated with LGRBs through both temporal and spatial coincidence, as well as spectroscopic confirmation (see e.g. Galama et al., 1998; Hjorth et al., 2003; Pian and Mazzali, 2006; Mirabal et al., 2006; Becerra et al., 2017). These SNe usually present peak absolute magnitudes of $M_{V}\approx-19$ to 20-20 around 10 to 15 days after the burst in the rest frame (e.g. Cano et al., 2017). This timescale is determined by the radioactive decay of 56Ni and 56Co.
At the luminosity distance of $d=4.9$ Gpc, standard type Ibc supernovae (SN 1998bw-like; Galama et al., 1999) would exhibit apparent peak magnitudes in the range of $\sim~24.5.0-25.5$, making them several orders of magnitude fainter than the late-time component observed in GRB 250221A. On the other hand, a very luminous supernova associated with a magnetar such as SN2011kl could be brighter by about one magnitude Greiner et al. (2015); Kann et al. (2019) and therefore comparable with the values exhibited for GRB 250221A in the rebrightening.
@@ 5.4 Gravitational Microlensing
It is tempting to associate the rebrightening with a microlensing event, assuming that G2 is the host and G1 is a galaxy on the line-of-sight. In the following, we briefly describe the formalism for GRB microlensing first introduced by Loeb and Perna (1998). As discussed in section 4.2, the temporal evolution of the spherical blast wave is described by the Blandford and McKee (1976) self-similar solution, with $\Gamma\propto R^{-3/2}$, after it has been decelerated by its interaction with the ISM.
The afterglow image on the plane of the sky is predicted to appear as a ring with a narrow fractional width, $W$, that depends on the observed frequency and whose radius grows with apparent time $T$ as a power law, with $R_{s}(T)\propto R/\Gamma\propto R^{5/2}\propto T^{5/8}$ where $R\propto\Gamma^{2}T\propto T^{1/4}$. In the optical band, $W$ is predicted to be of the order of a few percent (Loeb and Perna, 1998; Granot et al., 1999). After about a day after the trigger, the radius of the ring, $R_{s}(T)\approx 1.1\times 10^{17}~E_{\rm k,iso,52}^{1/8}~n_{1}^{-1/8}T_{\rm day}^{5/8}~{\rm cm}$ (Piran, 1999), becomes comparable to the Einstein radius of a solar mass lens, $r_{E}=\sqrt{(4GM/c^{2})(D_{L}D_{LS}/D_{S})}$ where $D_{L}$, $D_{LS}$, and $D_{S}$ are the lens-observer, lens-source and source-observer angular distances, respectively, $M$ is the mass of the lens.
The magnification of a uniform ring of fractional width $W$ and impact factor $b$ is given by Loeb and Perna 1998,
@@
Therefore, assuming that the lens is in G1 and the GRB is in G2, we get $D_{L}\approx 1037$ Mpc, $D_{LS}\approx 783$ Mpc, and $D_{S}\approx 1571$ Mpc, respectively (Planck Collaboration et al., 2020) and hence, $r_{E}\approx 3.1\times 10^{16}(M/M_{\odot})^{1/2}~{\rm cm}$. Since the radius of the ring depends on the kinetic energy and the density of the medium, we adopt here a range of values around the values shown in Figure 8. We find that for the following ranges $1\lesssim n_{1}\lesssim 5$ and $0.1\lesssim E_{\rm k,iso,52}\lesssim 10$, that are broadly consistent with those inferred in Sec. 4.2, the timescale for the excess demands $b<1$ and $M\sim 50~M_{\odot}$. The width of the excess demands a fractional width of the ring $W\sim 0.1$. However, this modelling does not allow accounting for the amplitude of the excess, as can be seen in Figure 9. We note that these microlensing calculations are done considering a uniform ring for the afterglow image (Loeb and Perna, 1998).
@@ 5.5 ISM Local Density Enhancement
As we describe in Section  4.2, the fireball model describes the afterglow emission as a smooth light curve that can be modelled with power-law segments. However, in reality, light curve variability is observed in several GRBs (Nakar et al., 2003; de Ugarte Postigo et al., 2005), and can be attributed to an increase in external density that the fireball interacts with, experiencing additional deceleration (Nakar and Granot, 2007).
@@
In contrast, for SGRBs, and likely GRB 250221A, this explanation does not apply. Their compact-object progenitors merge long after both stars have exploded as supernovae, leaving behind a CBM free of stellar-wind disturbances. Thus, any external density variations must stem from inhomogeneities in the ISM (Nakar and Granot, 2007).
Nevertheless, Nakar and Granot (2007) demonstrated that density fluctuations in the external medium have a minimal impact on the forward shock emission as the dynamics of the relativistic blast wave tend to suppress any resulting re-brightening. Furthermore, van Eerten et al. (2009) performed hydrodynamical simulations of a relativistic blast wave interacting with a wind termination shock, finding no significant variability in the optical light curve, even in the presence of large density contrasts. Therefore, these results favour scenarios involving refreshed shocks or multiple ejection episodes as more plausible explanations for the observed variability.
@@ 5.6.1 Reverse Shock Emission from Two-Shell Collision
To explain the rebrightening of the afterglow emission around starting from $T_{0}\sim 1$ day (hereafter $T_{\rm bright}$) , we first consider a scenario where a second (inner) shell of mass emitted by the central engine, with some delay of $t_{\rm em}$ in the engine frame, catches up from behind with the decelerating first (outer) shell and refreshes the shock by injecting more energy (e.g. Kumar and Piran, 2000; Sari and Mészáros, 2000; Zhang and Mészáros, 2002; Vlasis et al., 2011) (see Figure 10). We develop the model of two spherical shell collision in Appendix (A), where we calculate characteristic radii, timescales, and dynamics of the system (also see, e.g., Anderson et al., 2025). As the two shells collide (see Figure 10) a forward shock (FS) propagates into the outer shell and a reverse shock (RS) propagates into the inner shell. The shock jump conditions at the two shocks are different, as the outer shell is already relativistically hot while the inner shell is cold. The main parameters in the problem include the ratio of the kinetic energies of the two shells ($E_{(6),\rm k,iso}/E_{(3),\rm k,iso}$), the ratio of their initial bulk Lorentz factors ($\Gamma_{(3),0}/\Gamma_{(6),0}$), and the shock microphysical parameters at the two shocks. The parenthesized numbers in the subscript refer to the regions shown in Fig. 10.
Figure 10: Sketch (not to scale) of the different shocked and unshocked regions that arise in the collision of two kinetic-energy-dominated shells. The thick solid line indicates the pressure, or the internal energy density, in each region. The thin solid line indicates the comoving mass density. Figure adapted from Zhang and Mészáros (2002).
In the case of a binary merger, the two shells must be emitted with negligible delay, such that $t_{\rm em}\ll T_{\rm 90}$, and therefore the inner shell must have $\Gamma_{(6),0}<\Gamma_{(3),0}$. This condition also guarantees that the inner shell must be spreading radially before the collision, i.e. it is in the thin-shell regime. Also, in general, if $t_{\rm em}\ll T_{\rm bright}$, then the ratio of the initial bulk Lorentz factors is fixed (see Equation 16), such that
$\frac{\Gamma_{(3),0}}{\Gamma_{(6),0}}$ $\approx\frac{[4(T_{\rm bright}/T_{\rm dec})]^{3/8}}{2}$
@@
Since $\Gamma_{(3),0}\sim 400$ is obtained from the MCMC fits, this yields $\Gamma_{(6),0}\sim 8$. Figure 13 shows the dynamical evolution of the different shocked and unshocked regions, where we find that the RS remains non-relativistic as it crosses the inner shell, as expected for a thin shell.
As shown in Zhang and Mészáros (2002), the RS emission typically dominates over that coming from the other two shocked regions before it declines rapidly. The shock jump conditions in Equation 19 yield the energy and number density of the shocked medium, which we then use to calculate the characteristic synchrotron frequencies and the peak flux (e.g. Sari et al., 1998; Gill and Granot, 2018). In calculating the observed emission, we simplify to obtain radiation emitted only along the line-of-sight (LOS) without performing equal-arrival-time-surface (EATS) integration, with photon arrival time given by $T=(1+z)R/4\Gamma^{2}c$.
@@
The two shell collision scenario, as described here, demands a significantly large amount of energy to be residing in the inner shell. In this case, we find that $E_{(6),\rm k,iso}/E_{(3),\rm k,iso}\simeq 20$ is required in order to obtain the correct radius at which the two colliding shells are shocked (see Fig. 13) and produce the rebrightening, which for $\Gamma_{(6),0}\sim 8$ (see above) implies a large density in the inner shell compared to the outer one. Since the RS is non-relativistic, large densities are needed to obtain a bright emission. When such a large amount of energy is injected into the blast wave, it should also lead to very bright FS emission (not calculated in this model) after the RS emission fades, which would significantly overproduce the emission during the decaying phase of the rebrightening episode. Therefore, this model is disfavoured by the observations.
@@ 5.6.2 Energy Injection into the Forward Shock of a Jet
Next we consider an alternative model with continuous energy injection into the blast wave produced by a jet. Two different ways of doing so have been discussed in the literature for a spherical flow, where (i) the ejecta comprises a radial velocity stratification, with progressively slower moving inner shells trailing behind the faster outer shell (Rees and Mészáros, 1998; Sari and Mészáros, 2000), or (ii) a rapidly spinning central engine, e.g. a millisecond magnetar, continuously injects energy into the blast wave via a magneto-hydrodynamical (MHD) wind as it spins down (Dai and Lu, 1998b, a; Zhang and Mészáros, 2001, 2002). In both scenarios, the afterglow light curve shows an achromatic bump or rebrightening as the energy is injected, with the post-injection light curve following the same temporal decay as that of the pre-injection one in the absence of spectral break passage. Both scenarios have been used in many earlier works (e.g., Laskar et al., 2018; Schroeder et al., 2024; de Wet et al., 2024; Schroeder et al., 2025a) to explain rebrightening features in afterglow light curves, with the assumption that the trailing ejecta in model (i) only catches up with the faster moving blast wave at the time of the rebrightening.
@@
so that the energy of the blast wave grows as $E(R)\propto R^{1+\ell}$. From energy conservation in the blast wave (Blandford and McKee, 1976; Chiang and Dermer, 1999; Huang et al., 1999; Zhang and Mészáros, 2002), we get
@@
To solve for the dynamical evolution of the spherical blast wave with energy injection, we use the numerical code of Gill and Granot (2023) that produces EATS-integrated afterglow emission. Figure 12 shows the afterglow light curve where we take the best-fit solution from Figure 8 and inject $E_{\rm inj}\simeq 4.5E_{\rm k,iso}$ into the blast wave with $\ell=0.1$. During the energy injection phase, the light curves for the radio ($\nu_{m}<\nu_{R}<\nu_{a}<\nu_{c}$), optical ($\nu_{a}<\nu_{O}<\nu_{c}$), and X-rays ($\nu_{c}<\nu_{X}$) show distinct temporal trends since all three are located on different power-law segments of the synchrotron spectrum. Here $\nu_{m}$ and $\nu_{c}$ are the characteristic injection and cooling-break frequencies, as defined earlier, and $\nu_{a}$ is the self-absorption frequency. Since the radio emission is self-absorbed, the rebrightening shows a temporal lag of $\Delta T\sim 4$ days with respect to that seen in optical and X-rays.
To explain the late-time radio and X-ray emission, a jet break in the light curve at $T\sim 10^{5}$ s is required, with a jet half-opening angle of $\theta_{j}=11.5^{\circ}$. We introduce the jet break in our spherical model by manually integrating the emission over the surface of the flow up to the sharp edge of the jet at $\theta_{j}$, for the same parameters obtained from the MCMC fit of the spherical model. An on-axis ($\theta_{\rm obs}=0$) observer would see a jet break in the afterglow light curve when the angular size of the beaming cone around the line-of-sight exceeds the opening angle of the jet, i.e. when $1/\Gamma=\theta_{j}$, which would occur at the time $T_{j}=(\Gamma_{0}\theta_{j})^{8/3}T_{\rm dec}\simeq 2\times 10^{5}$ s for $\Gamma_{0}=400$ and $T_{\rm dec}\simeq 2$ s. The true jet kinetic energy in this scenario can be estimated from the total isotropic-equivalent energy, with $E_{k}\simeq(\theta_{j}^{2}/2)E_{\rm k,iso,tot}\simeq 4.4\times 10^{51}$ erg where $E_{\rm k,iso,tot}=\mathbf{E_{\rm inj}+E_{\rm k,iso}}\simeq 5.5E_{\rm k,iso}$, and $E_{\rm k,iso}$ is the energy of the fastest moving material.
@@
In the first scenario, where the ejecta has a radial gradient in velocity, so that the mass above a given bulk-$\Gamma$ scales as $M(>\Gamma)\propto\Gamma^{-s}$ (Rees and Mészáros, 1998; Sari and Mészáros, 2000), the energy in the blast wave grows as $E(>\Gamma)\propto\Gamma^{1-s}\propto T^{-3(1-s)\over(7+s)}$ after the fastest moving material decelerates. The scaling of the blast wave energy with $T$ in this model corresponds to $s=(4+\ell)/(2-\ell)\simeq 2.2$. Since $s>1$, both the mass and energy in the ejecta are dominated by the slower moving material. In the second scenario, energy is injected by the spinning down magnetar with power $L(T)\propto T^{q}$, where the blast wave energy grows as $E\propto T^{1+q}$ (Zhang and Mészáros, 2001). Comparison of this scaling with our model gives $q=2(\ell-1)/(3-\ell)=-0.62$. When q>1q>-1 the injected energy is expected to dominate the blast wave dynamics over the impulsive, i.e. without energy injection, blast wave as the injected kinetic energy will exceed that of the impulsive shell (see above). In this scenario, the injected power is produced by electromagnetic losses due to dipolar spindown of the magnetar, and this power is expected to follow the scaling with observer-frame time, with $L(T)\propto T^{0}$ for $T\ll T_{\rm sd}$ and $L(T)\propto T^{-2}$ for $T\gg T_{\rm sd}$. The characteristic spin-down time is given by $T_{\rm sd}=3c^{3}I/B_{p}^{2}R_{\rm NS}^{6}\Omega_{0}^{2}=2.05\times 10^{3}I_{45}B_{p,15}^{-2}P_{0,-3}^{2}R_{\rm NS,6}^{-6}$ s, where $B_{p}$ is the surface polar magnetic field, $P_{0}$ is the initial spin period, and $I$ and $R_{\rm NS}$ are the moment of inertia and radius of the neutron star. The obtained value of q=0.62q=-0.62 does not match the expected asymptotic power-law slopes of the injected power both before and after $T_{\rm sd}$. Therefore, this simple description of the magnetar scenario is not strongly favoured, however, a more detailed MCMC fit is required to rule it out.